OpenStudy (anonymous):

How many possible solutions are there for these equations? x - 3y = -3, -2x + 6y = 12

OpenStudy (texaschic101):

x - 3y = -3 (put in y = mx + b form) -3y = -x - 3 -- divide both sides by -3 y = 1/3x + 1 -2x + 6y = 12 (put in y = mx + b form) 6y = 2x + 12 -- divide both sides by 6 y = 1/3x + 2 okay...learn this if the slopes (numbers in the m position) are the same, and the y intercept (numbers in the b position) are the same, then it is the same line with infinite solutions. if the slopes are the same and the y intercepts different, then it is a parallel line with no solutions because the lines never intersect. if the slopes are different and the y intercepts different, then it just has one solution conclusion : slopes are the same, y intercepts are different..making the lines parallel with no solutions